# Qomhrá: A Bilingual Irish and English Large Language Model

### v2

@@ 1 Introduction
Whilst progress has been made in speech synthesis and recognition for Irish lonergan-etal-2022-automatic, Irish remains the least-supported official European language in terms of language technology ele2022irish. To address this, we present Qomhrá, an open-weights bilingual Irish and English LLM, developed under extremely low-resource constraints. Qomhrá supports the subsidiary development of chatbots and other NLP tools, applicable across education, translation, and public services, enabling access to language technologies for the Irish-language community. Crucially, Qomhrá fosters technological sovereignty for the Irish language by reducing reliance on proprietary API-based models. This mitigates risks associated with cost volatility and data privacy, while ensuring the community retains control over the adaptation of models to specific use cases.
@@
Definitions of low-resource vary from the socio-political to human expertise and data availability nigatu-etal-2024-zenos. In the context of LLM development, we classify Irish as a low-resource language due to data scarcity, or, more precisely, the lack of permissively-licensed labelled data for instruction tuning and alignment. This data requires explicit human annotation, making it particularly valuable for training and evaluating LLMs but also it is typically costly to create.
@@
An L2 speaker is considered in this context because they drastically outnumber L1 speakers of Irish. As of the 2022 census of Ireland cso2022-education-irish, 40% of people indicated that they speak Irish, of which, only 10% indicated that they speak “very well”. Therefore, if L2 speakers prove effective annotators, the cost of annotation can be reduced drastically. The same reasoning is applied for the inclusion of LLMs for annotation.
@@
    The release of Qomhrá, an 8B parameter open-weight Irish and English LLM  [^1] Anonymised for submission, including a language-aware quantized version  [^2] Anonymized for submission.
@@
    A 30K sample parallel English-Irish instruction tuning dataset that significantly improves the LLM’s ability to follow instructions  [^3] Anonymised for submission.
@@
    We release a 1K sample Irish human preference dataset that is shown to align with an L1 Irish speaker  [^4] Anonymised for submission.
@@ 2.1 LLMs for Low-Resource Languages
While major languages, particularly English, dominate interest in state-of-the-art LLM performance, there have also been efforts that have focused on adapting models to low-resource languages. Initial efforts for Irish focused on encoder-only models such as gaBERT and gaELECTRA barry-etal-2022-gaBERT. More recently, the focus has shifted to generative decoder models, such as UCCIX for Irish tran2024uccixirishexcellencelargelanguage and elsewhere, Latxa for Basque etxaniz-etal-2024-latxa, adapting to their respective low-resource languages via continued pre-training gururangan-etal-2020-CPT. We note that both efforts report cases of catastrophic forgetting cata_MCCLOSKEY1989109, where English performance declines, and thus we include a significantly higher proportion of English training data for bilingual capabilities.
@@ 2.2 Instruction Tuning
For high-resource languages, instructions annotated by humans are abundant. In low-resource settings, they are often either low-quality or absent. A prevailing strategy is translate-train conneau-etal-2018-xnli, where data is translated from source to target language and then used to train the model. This is effective at augmenting low-resource instruction tuning data singh2024ayadatasetopenaccesscollection.
We adopt this methodology by translating the Dolly V2 dataset DatabricksBlog2023DollyV2, but we extend the pipeline by evaluating the translator model first, since in low-resource settings model providers often do not evaluate low-resource languages. This extension helps others looking to select models for Irish data synthesis and highlights a key step for others seeking to synthesise data in other low-resource languages.
@@ 2.3 Alignment
After instruction following, the next step is to align models to human preferences, which has typically relied on Reinforcement Learning from Human Feedback (RLHF) Christiano_1017_RLHF, a procedure which is computationally complex and costly. Direct Preference Optimization (DPO) DPO_Rafilov_2023 offers a more stable, data-efficient alternative but requires paired “chosen” and “rejected” response data, which does not exist for Irish. Synthetic data generation for DPO has been explored for other languages kazakh-pseudo-DPO, often using a teacher model to score responses. We propose a simplified, novel approach that explicitly prompts a strong multilingual teacher (Gemini-2.5-Pro) to synthesise the preference signal itself by generating contrasting translation qualities, thereby bypassing the need for a fine-tuned teacher model.
@@ 3.1 Continued Pre-training Data
As with the GaBERT and UCCIX models, we use the available web-crawled Irish data, drawing from the de-duplicated dataset released by the UCCIX team tran2024uccixirishexcellencelargelanguage. Additionally, a subset of the National Corpus of Irish was made available for this project under the CC BY-SA 4.0 license. Due to copyright restrictions, the corpus was shuffled at the sentence level and copyright-protected data was not shared. An overview of the pre-training data and sources is shown in Table 1.
@@ 3.1.1 De-duplication
Near duplicate documents introduce redundancy in training. In the context of this project, the data released by the UCCIX team has already undergone pre-processing to remove duplicates. Therefore, our most pertinent question was whether the CNG and Bible data were already contained containment_broder in the UCCIX data. Containment (MinHash) is defined below in Equation 1, where $|A|<|B|$. To measure the containment, the text was lower-cased and punctuation was removed. Documents were transformed (L) into sets of unique 5-gram deterministic hashes in line with UCCIX pre-processing methodology. tran2024uccixirishexcellencelargelanguage.
@@ 3.1.2 Segmentation
The end of document token <|endoftext|> QwenConcepts was inserted between samples across all data sources, preventing the model from inferring dependencies between unrelated documents and leveraging a special token it is already trained with. For the Dáil dataset, this corresponded to between speaker utterances; for CNG, the end of text token was appended to each sentence (given the sentence-level data shuffling of this dataset); and for the UCCIX dataset between each textual sample in the dataset. Each parallel eng-gle sample was explicitly prepended with its respective language tag: "[eng]"/"[ga]".
@@ 3.2 Continued Pre-training
While methods like in-context learning ICL_cahyawijaya-etal-2024-llms and parameter efficient fine-tuning (PEFT) LoRA_2021; QLoRA offer efficient adaptation, continued pre-training (CPT) gururangan-etal-2020-CPT is better suited for the large domain shift of a new language Lu_Luu_Buehler_2025_CPT_LORA.
@@ 3.2.1 Training Configuration and Hyper-parameters
A per-device batch size of one with gradient accumulation ($\times 8$) was used to stay within memory constraints. DeepSpeed ZeRO-2 enabled data parallelism with both optimiser and gradient partitioning across GPUs. Gradient checkpointing reduced memory overhead. Training ran for two epochs to balance convergence with constraints, with validation monitoring from Weights & Biases wandb and a test script to load model from checkpoint and run test generation every 3K steps to assess progress.
@@
The BF16 google-cloud-bfloat16 floating-point format was used as it approximates the dynamic range of FP32 with the 50% memory reduction FP16 provides kalamkar2019stud_BF16y, which is important to maximise compute resources while maintaining training stability.
@@ 3.3 Instruction Tuning
Instruction tuning transitions an LLM from a token predictor to a chatbot, useful for interfacing with humans. Qomhrá is trained on instruction-response pairs which requires significantly less data wei2022FLAN. We translate the Dolly V2 DatabricksBlog2023DollyV2 dataset to Irish. A translation of this dataset contributes a parallel Irish and English instruction-tuning dataset. The quality, however, is bound by the machine translation model. As such, an experiment was set up to determine the strongest closed-weight LLM for Irish, outlined in Section 4.1.
@@ 3.3.1 Training Configuration and Hyper-parameters
The Qwen chat template was applied to prompts. The thinking tags were removed as the goal was not to develop a reasoning model. BF16 format was used and learning rate hyperparameters were tuned in line with the Unsloth LoRA Hyper-parameter Guide unsloth2025lora. The learning rate was $2e^{-4}$ with AdamW optimiser, rank 16, $\alpha=32$, weight decay =0.01=0.01 and 3% warm up ratio for fast adaptation.
@@ 3.4 Human Feedback
To do this, Gemini-2.5-Pro was instructed to translate each English instruction response sample, where the prompt is identical but the response varies, where an L1 speaker is expected to “accept” one response and “reject” the other. The LIMA dataset LIMA was selected for translation, due to its effectiveness at small scale. The top-ranked model, Gemini-2.5-Pro, was used to synthesise the preference dataset. The following is the prompt we designed, specifying high contrast and strict JSON formatting for robust API parsing.
@@ 4.1 Translator Selection Methodology
Each model was provided with Irish language text from the Irish parliament (Dáil) and Wikipedia and asked to generate a prompt-response pair. This follows methodology applied to generate synthetic Kazakh instruction-tuning data laiyk-etal-2025-instruction.
120 annotations were completed by both the L2 and the L1 speaker, whereas the LLMs annotated all 600 samples. The ranking model selected was the Bradley-Terry model due to its strong performance dealing with small sample sizes daynauth-etal-2025-ranking.
@@ 4.3.1 Benchmark Descriptions
Cloze-gle tran2024uccixirishexcellencelargelanguage evaluates the model’s grasp of Irish grammatical gender and person agreement (masculine/feminine/plural). The model is presented with three candidate sentences differing only by a single pronoun (e.g., é, í, iad). The log-likelihood (LL) of each full sentence is computed. Since the candidates differ by only one word, the sentence with the highest likelihood represents the model’s prediction. In Table 5, the noun faoiseamh is masculine. The model assigns the highest likelihood (-38.2) to the sentence containing the masculine pronoun é, matching the target.
@@
SIB-gle adelani2024sib200simpleinclusivebig tests topic modelling capabilities. Unlike the zero-shot Cloze task, this is a 10-shot evaluation. The model is presented with input text and must assign it to one of seven high-level topics (e.g., eolaíocht/teicneolaíocht [science/tech], polaitíocht [politics]). The model scores each topic label as a continuation; the label with the highest probability is selected. In Table 6, the text describes how a nuclear bomb is made and the model correctly assigns the highest probability to the science category. Only 2 of the 7 categories are displayed in Table 6 for brevity.
@@
IQA-gle & IQA-eng tran2024uccixirishexcellencelargelanguage assess multiple-choice question answering in both English and Irish using a 5-shot context. For a given question, the model is presented with four candidate answers. The log-likelihood of each candidate answer is calculated given the question context. Table 7 illustrates an example of IQA-eng where the model correctly identifies that Irish is taught as a compulsory subject. The method is the same for IQA-gle.
@@
BLEU (gle $\leftrightarrow$ eng) lankford-etal-2022-gahealth evaluates translation quality via BLEU-4 in a 5-shot setting on health data. BLEU-4 measures the n-gram overlap (up to 4-grams) between the model’s response and a gold-standard reference. Table 8 shows a sample generation where Qomhrá’s response omits the definitive plural article na, this difference is highlighted in bold.
@@
NQ-eng nq-2019 evaluates world knowledge via open-ended generation (5-shot). Success is measured by an exact match: the generated text must be a sub-string of a valid reference answer. Table 9 demonstrates a failure case. Although the model correctly identifies the date ("October 24"), it fails to specify the year required by the reference. This strict metric explains the lower scores observed in Table 4.
@@ 4.3.2 Benchmark Results
Qomhrá outperformed UCCIX by 29.2% in Irish as opposed to 44.4% in English on the IQA benchmark. The information from these questions is the same so language capability must be the factor that caused Qomhrá to outperform UCCIX, indicating improved English performance retention for Qomhrá, likely due to its higher proportion of English data at 25% compared to UCCIX’s 1% . Qomhrá did not improve performance with the second epoch of training. This is in line with expectations that base models saturate after one epoch of CPT Lu_Luu_Buehler_2025_CPT_LORA. Therefore in future development of the model, CPT would only be trained for one epoch to prevent redundant compute expenditure.
@@ 4.4 Instruction Tuning
This improved performance after instruction tuning was unsurprising as its intention is to improve unseen task performance wei2022FLAN. Future work could extend instruction tuning by adding cultural knowledge through the translation of parliamentary and government data as seen to be effective for Kazakh laiyk-etal-2025-instruction.
@@ 5 Conclusion
LLM CPT requires significant compute power which contributes to greenhouse gas emissions, as such it is important to quantify our impact. The power usage of an Nvidia H100 GPU is 700W. We trained for 44.196 hours on 2 GPUs. The carbon intensity was approximately 200 gCO2/kWh CI-Eirgrid, which gives a total of 12.37kg CO2. This can be understood relative to the average of 10.4 tonnes of CO2 per capita in Ireland CSO-Env. Efficient corpus sampling methods can be used in future work for more efficient training.

### v3

@@ 1 Introduction
Whilst progress has been made in speech synthesis and recognition for Irish Lonergan et al. (2022), Irish remains the least-supported official European language in terms of language technology Lynn (2022). To address this, we present Qomhrá, an open-weights bilingual Irish and English LLM, developed under extremely low-resource constraints. Qomhrá supports the subsidiary development of chatbots and other NLP tools, applicable across education, translation, and public services, enabling access to language technologies for the Irish-language community. Crucially, Qomhrá fosters technological sovereignty for the Irish language by reducing reliance on proprietary API-based models. This mitigates risks associated with cost volatility and data privacy, while ensuring the community retains control over the adaptation of models to specific use cases.
@@
Definitions of low-resource vary from the socio-political to human expertise and data availability Nigatu et al. (2024). In the context of LLM development, we classify Irish as a low-resource language due to data scarcity, or, more precisely, the lack of permissively-licensed labelled data for instruction tuning and alignment. This data requires explicit human annotation, making it particularly valuable for training and evaluating LLMs but also it is typically costly to create.
@@
An L2 speaker is considered in this context because they drastically outnumber L1 speakers of Irish. As of the 2022 census of Ireland Central Statistics Office (2022), 40% of people indicated that they speak Irish, of which, only 10% indicated that they speak “very well”. Therefore, if L2 speakers prove effective annotators, the cost of annotation can be reduced drastically. The same reasoning is applied for the inclusion of LLMs for annotation.
@@
    The release of Qomhrá, an 8B parameter open-weight Irish and English LLM 111huggingface.co/jmcinern/Qomhra, including a language-aware quantized version 222huggingface.co/jmcinern/Qomhra-AWQ.
@@
    A 30K sample parallel English-Irish instruction tuning dataset that significantly improves the LLM’s ability to follow instructions 333huggingface.co/datasets/jmcinern/Dolly-V2-gle.
@@
    We release a 1K sample Irish human preference dataset that is shown to align with an L1 Irish speaker 444huggingface.co/datasets/jmcinern/LIMA-gle-DPO.
@@ 2.1 LLMs for Low-Resource Languages
While major languages, particularly English, dominate interest in state-of-the-art LLM performance, there have also been efforts that have focused on adapting models to low-resource languages. Initial efforts for Irish focused on encoder-only models such as gaBERT and gaELECTRA Barry et al. (2022). More recently, the focus has shifted to generative decoder models, such as UCCIX for Irish Tran et al. (2024) and elsewhere, Latxa for Basque Etxaniz et al. (2024), adapting to their respective low-resource languages via continued pre-training Gururangan et al. (2020). We note that both efforts report cases of catastrophic forgetting McCloskey and Cohen (1989), where English performance declines, and thus we include a significantly higher proportion of English training data for bilingual capabilities.
@@ 2.2 Instruction Tuning
For high-resource languages, instructions annotated by humans are abundant. In low-resource settings, they are often either low-quality or absent. A prevailing strategy is translate-train Conneau et al. (2018), where data is translated from source to target language and then used to train the model. This is effective at augmenting low-resource instruction tuning data Singh et al. (2024).
We adopt this methodology by translating the Dolly V2 dataset Conover et al. (2023), but we extend the pipeline by evaluating the translator model first, since in low-resource settings model providers often do not evaluate low-resource languages. This extension helps others looking to select models for Irish data synthesis and highlights a key step for others seeking to synthesise data in other low-resource languages.
@@ 2.3 Alignment
After instruction following, the next step is to align models to human preferences, which has typically relied on Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017), a procedure which is computationally complex and costly. Direct Preference Optimization (DPO) Rafailov et al. (2023) offers a more stable, data-efficient alternative but requires paired “chosen” and “rejected” response data, which does not exist for Irish. Synthetic data generation for DPO has been explored for other languages Kadyrbek et al. (2025), often using a teacher model to score responses. We propose a simplified, novel approach that explicitly prompts a strong multilingual teacher (Gemini-2.5-Pro) to synthesise the preference signal itself by generating contrasting translation qualities, thereby bypassing the need for a fine-tuned teacher model.
@@ 3.1 Continued Pre-training Data
As with the GaBERT and UCCIX models, we use the available web-crawled Irish data, drawing from the de-duplicated dataset released by the UCCIX team Tran et al. (2024). Additionally, a subset of the National Corpus of Irish was made available for this project under the CC BY-SA 4.0 license. Due to copyright restrictions, the corpus was shuffled at the sentence level and copyright-protected data was not shared. An overview of the pre-training data and sources is shown in Table 1.
@@ 3.1.1 De-duplication
Near duplicate documents introduce redundancy in training. In the context of this project, the data released by the UCCIX team has already undergone pre-processing to remove duplicates. Therefore, our most pertinent question was whether the CNG and Bible data were already contained Broder (1997) in the UCCIX data. Containment (MinHash) is defined below in Equation 1, where $|A|<|B|$. To measure the containment, the text was lower-cased and punctuation was removed. Documents were transformed (L) into sets of unique 5-gram deterministic hashes in line with UCCIX pre-processing methodology. Tran et al. (2024).
@@ 3.1.2 Segmentation
The end of document token <|endoftext|> Qwen Team (2024) was inserted between samples across all data sources, preventing the model from inferring dependencies between unrelated documents and leveraging a special token it is already trained with. For the Dáil dataset, this corresponded to between speaker utterances; for CNG, the end of text token was appended to each sentence (given the sentence-level data shuffling of this dataset); and for the UCCIX dataset between each textual sample in the dataset. Each parallel eng-gle sample was explicitly prepended with its respective language tag: "[eng]"/"[ga]".
@@ 3.2 Continued Pre-training
While methods like in-context learning Cahyawijaya et al. (2024) and parameter efficient fine-tuning (PEFT) Hu et al. (2021); Dettmers et al. (2023) offer efficient adaptation, continued pre-training (CPT) Gururangan et al. (2020) is better suited for the large domain shift of a new language Lu et al. (2025).
@@ 3.2.1 Training Configuration and Hyper-parameters
A per-device batch size of one with gradient accumulation ($\times 8$) was used to stay within memory constraints. DeepSpeed ZeRO-2 enabled data parallelism with both optimiser and gradient partitioning across GPUs. Gradient checkpointing reduced memory overhead. Training ran for two epochs to balance convergence with constraints, with validation monitoring from Weights & Biases Biewald (2020) and a test script to load model from checkpoint and run test generation every 3K steps to assess progress.
@@
The BF16 Google Cloud (2019) floating-point format was used as it approximates the dynamic range of FP32 with the 50% memory reduction FP16 provides Kalamkar et al. (2019), which is important to maximise compute resources while maintaining training stability.
@@ 3.3 Instruction Tuning
Instruction tuning transitions an LLM from a token predictor to a chatbot, useful for interfacing with humans. Qomhrá is trained on instruction-response pairs which requires significantly less data Wei et al. (2022). We translate the Dolly V2 Conover et al. (2023) dataset to Irish. A translation of this dataset contributes a parallel Irish and English instruction-tuning dataset. The quality, however, is bound by the machine translation model. As such, an experiment was set up to determine the strongest closed-weight LLM for Irish, outlined in Section 4.1.
@@ 3.3.1 Training Configuration and Hyper-parameters
The Qwen chat template was applied to prompts. The thinking tags were removed as the goal was not to develop a reasoning model. BF16 format was used and learning rate hyperparameters were tuned in line with the Unsloth LoRA Hyper-parameter Guide Unsloth Documentation (2025). The learning rate was $2e^{-4}$ with AdamW optimiser, rank 16, $\alpha=32$, weight decay =0.01=0.01 and 3% warm up ratio for fast adaptation.
@@ 3.4 Human Feedback
To do this, Gemini-2.5-Pro was instructed to translate each English instruction response sample, where the prompt is identical but the response varies, where an L1 speaker is expected to “accept” one response and “reject” the other. The LIMA dataset Zhou et al. (2023) was selected for translation, due to its effectiveness at small scale. The top-ranked model, Gemini-2.5-Pro, was used to synthesise the preference dataset. The following is the prompt we designed, specifying high contrast and strict JSON formatting for robust API parsing.
@@ 4.1 Translator Selection Methodology
Each model was provided with Irish language text from the Irish parliament (Dáil) and Wikipedia and asked to generate a prompt-response pair. This follows methodology applied to generate synthetic Kazakh instruction-tuning data Laiyk et al. (2025).
120 annotations were completed by both the L2 and the L1 speaker, whereas the LLMs annotated all 600 samples. The ranking model selected was the Bradley-Terry model due to its strong performance dealing with small sample sizes Daynauth et al. (2025).
@@ 4.3.1 Benchmark Descriptions
Cloze-gle Tran et al. (2024) evaluates the model’s grasp of Irish grammatical gender and person agreement (masculine/feminine/plural). The model is presented with three candidate sentences differing only by a single pronoun (e.g., é, í, iad). The log-likelihood (LL) of each full sentence is computed. Since the candidates differ by only one word, the sentence with the highest likelihood represents the model’s prediction. In Table 5, the noun faoiseamh is masculine. The model assigns the highest likelihood (-38.2) to the sentence containing the masculine pronoun é, matching the target.
@@
SIB-gle Adelani et al. (2024) tests topic modelling capabilities. Unlike the zero-shot Cloze task, this is a 10-shot evaluation. The model is presented with input text and must assign it to one of seven high-level topics (e.g., eolaíocht/teicneolaíocht [science/tech], polaitíocht [politics]). The model scores each topic label as a continuation; the label with the highest probability is selected. In Table 6, the text describes how a nuclear bomb is made and the model correctly assigns the highest probability to the science category. Only 2 of the 7 categories are displayed in Table 6 for brevity.
@@
IQA-gle & IQA-eng Tran et al. (2024) assess multiple-choice question answering in both English and Irish using a 5-shot context. For a given question, the model is presented with four candidate answers. The log-likelihood of each candidate answer is calculated given the question context. Table 7 illustrates an example of IQA-eng where the model correctly identifies that Irish is taught as a compulsory subject. The method is the same for IQA-gle.
@@
BLEU (gle $\leftrightarrow$ eng) Lankford et al. (2022) evaluates translation quality via BLEU-4 in a 5-shot setting on health data. BLEU-4 measures the n-gram overlap (up to 4-grams) between the model’s response and a gold-standard reference. Table 8 shows a sample generation where Qomhrá’s response omits the definitive plural article na, this difference is highlighted in bold.
@@
NQ-eng Kwiatkowski et al. (2019) evaluates world knowledge via open-ended generation (5-shot). Success is measured by an exact match: the generated text must be a sub-string of a valid reference answer. Table 9 demonstrates a failure case. Although the model correctly identifies the date ("October 24"), it fails to specify the year required by the reference. This strict metric explains the lower scores observed in Table 4.
@@ 4.3.2 Benchmark Results
Qomhrá outperformed UCCIX by 29.2% in Irish as opposed to 44.4% in English on the IQA benchmark. The information from these questions is the same so language capability must be the factor that caused Qomhrá to outperform UCCIX, indicating improved English performance retention for Qomhrá, likely due to its higher proportion of English data at 25% compared to UCCIX’s 1% . Qomhrá did not improve performance with the second epoch of training. This is in line with expectations that base models saturate after one epoch of CPT Lu et al. (2025). Therefore in future development of the model, CPT would only be trained for one epoch to prevent redundant compute expenditure.
@@ 4.4 Instruction Tuning
This improved performance after instruction tuning was unsurprising as its intention is to improve unseen task performance Wei et al. (2022). Future work could extend instruction tuning by adding cultural knowledge through the translation of parliamentary and government data as seen to be effective for Kazakh Laiyk et al. (2025).
@@ 5 Conclusion
LLM CPT requires significant compute power which contributes to greenhouse gas emissions, as such it is important to quantify our impact. The power usage of an Nvidia H100 GPU is 700W. We trained for 44.196 hours on 2 GPUs. The carbon intensity was approximately 200 gCO2/kWh EirGrid (2025), which gives a total of 12.37kg CO2. This can be understood relative to the average of 10.4 tonnes of CO2 per capita in Ireland Central Statistics Office of Ireland (2025). Efficient corpus sampling methods can be used in future work for more efficient training.